### v1

# Revisiting the Radial Metallicity Gradient-Age Relation in the Milky Way’s Thin and Thick Disks

Abstract
Galactic disks typically exhibit a negative radial metallicity gradient, indicating faster enrichment in the inner regions. Recent studies report that this gradient becomes flatter with increasing stellar age in the Milky Way’s (MW) thin disk, while the thick disk exhibits a mildly positive gradient across all ages. In this work, we revisit the metallicity gradient–age relation (MGAR) in both the thin and thick disks of the MW. We use spectroscopic data from LAMOST DR8 and stellar ages calibrated with asteroseismology. Our results show a steadily flattening MGAR in the thin disk and confirm a positive gradient $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$ in the thick disk. The flattening in the thin disk may be caused by large-scale radial migration induced by transient spiral arms, or by a time-dependent steepening of the interstellar medium (ISM) metallicity gradient as suggested by recent FIRE2 simulations. The positive gradient in the thick disk may reflect early enrichment of the outer regions by strong feedback or starburst-driven outflows in a turbulent, gas-rich proto-disk. These findings suggest distinct chemodynamical evolution paths for the MW’s thin and thick disks and provide valuable constraints for future models of Galactic chemical evolution.
Milky Way dynamics (1051) — Milky Way disk (1050) — Milky Way evolution (1052)
1 Introduction
A cornerstone of galaxy evolution is the negative radial metallicity gradient, which describes the decrease in stellar metallicity with increasing galactocentric radius. Matteucci & Francois (1989) reproduced this trend using “inside-out” formation models, in which the inner disk forms stars more rapidly and enriches earlier than the outer regions. This scenario naturally results in a negative metallicity gradient for stars formed from the interstellar medium (ISM, Matteucci, 2021). Following this framework, Minchev et al. (2018) found that the ISM gradient flattens over time as the outer disk becomes progressively enriched and approaches the central metallicity.
In this context, we refer to the stellar birth gradient as the radial metallicity gradient imprinted on stars at their formation, which inherits from the corresponding ISM gradient. Reconstructing the time sequence of these birth gradients would require direct knowledge of the metallicity distribution of newly formed stars at each epoch, which is not directly observable. Instead, what we can measure today is the metallicity gradient of mono-age stellar populations. By using stellar age as a proxy for look-back time, these present-day gradients give rise to the metallicity gradient–age relation (MGAR). It is also important to note that the present-day radial distribution of a mono-age stellar population does not necessarily preserve the distribution of their birth radii, as stars can migrate significantly over their lifetimes. Consequently, the observed MGAR is shaped by both the time evolution of the stellar birth gradient and the subsequent dynamical history of the Milky Way (MW).
The pioneering study on the MGAR in our MW by Mayor (1976) measured a mean radial metallicity gradient of $0.05\pm 0.01\,\mathrm{dex\,kpc^{-1}}$, which steepens to $0.10\pm 0.02\,\mathrm{dex\,kpc^{-1}}$ for younger stars. Using effective temperature ($T_{\mathrm{eff}}$) of main-sequence stars as a proxy for age, Yu et al. (2012) found that the gradient steepens with increasing $T_{\mathrm{eff}}$ (typically associated with decreasing age), but shows no trend for stars with $T_{\mathrm{eff}}<6,000\,\mathrm{K}$—those with main-sequence lifetimes $\gtrsim 10$ Gyr. These early findings hinted at a flattening gradient for increasingly older stars in the MW’s thin disk.
Large-scale spectroscopic surveys have greatly facilitated precise age estimates, enabling direct MGAR measurements (e.g., Vickers et al., 2021; Wang et al., 2023b; Ratcliffe et al., 2023). Vickers et al. (2021, hereafter V21) used LAMOST DR5 and reported a flattening gradient in the thin disk, starting at $0.075\,\mathrm{dex\,kpc^{-1}}$ for the youngest stars and flattening with a rate of $0.003\,\mathrm{dex\,kpc^{-1}\,Gyr^{-1}}$. They also found a slightly positive gradient of $\sim 0.02\,\mathrm{dex\,kpc^{-1}}$ for all ages in the thick disk, consistent with constantly positive gradients of old, cool main-sequence stars ($T_{\mathrm{eff}}<6,000\,\mathrm{K}$) in Yu et al. (2012), which are likely from the thick disk.
Radial migration has been proposed as a major process that alters stellar positions and flattens metallicity gradients (Sellwood & Binney, 2002). This mechanism, sometimes dubbed “churning”, causes significant radial mixing; non-axisymmetric perturbations such as transient spiral arms can induce large-scale stellar migration by exchanging stars across the corotation radius. In contrast, “blurring” refers to orbital heating that increases the amplitude of stellar epicycles without changing their guiding centers. Churning flattens the MGAR as older stars have had more time to migrate, resulting in a flatter radial metallicity distribution.
Recent simulations support the impact of radial migration. Lu et al. (2024) and Ratcliffe et al. (2023) analyzed the correlation between [Fe/H] spreads and stellar birth gradients in mono-age simulated populations, and recovered a time-dependent flattening of birth gradients. Anders et al. (2023) further compared observed MGAR from APOGEE DR17 with modeled ISM gradients evolution from Ratcliffe et al. (2023) to demonstrate the role of migration.
However, the migration interpretation has been challenged by recent FIRE2 simulations (Graf et al., 2024, 2025). They found that present-day MGARs closely follow the stellar birth gradient with minimal influence from churning in these simulations. As gas-phase metal gets mixed before star formation, the role of churning is significantly weakened.
Moreover, Graf et al. (2024) argued that vertical evolution—where turbulent gas settles over time, known as the “upside-down” scenario—may play a more dominant role than inside-out radial growth in shaping Galactic chemo-dynamical history. This suggests the need to consider vertical processes and possible differences in the evolutionary pathways of the thin and thick disks, as also done by V21.
To constrain the MGAR in the MW’s thin and thick disks, we follow an approach similar to V21. We isolate “pure” thin and thick disk populations using both geometric and chemical criteria, and compute MGAR separately for each population. Stellar kinematics and abundances are taken from LAMOST DR8 (Wang et al., 2022a), and stellar ages are calibrated from asteroseismic data provided by the Kepler mission (Wang et al., 2023a), offering an alternative to traditional isochrone fitting.
This paper is organized as follows. Section 2 describes the data selection, disk classification, orbital calculation, and gradient fitting method. Section 3 presents the MGAR measurements in both the thin and thick disks. Section 4 discusses the implications of our results and compares MGAR in the thin disk to the literature. Section 5 summarizes our conclusions.
2 Data and Method
The Large Area Multi-Object Spectroscopic Telescope (LAMOST), located at Xinglong Station of the National Astronomical Observatory of China, is a 4-meter Schmidt telescope with a 5° field of view. It is one of the most extensive stellar spectroscopic surveys and has released 13 data sets to date. In this work, we use LAMOST Data Release 8 (DR8; Wang et al., 2022a) combined with stellar ages from Wang et al. (2023a), which trained a neural network using asteroseismic data from the LAMOST-Kepler survey.
Our cross-matched sample contains 986,405 unique stars: 688,369 red giant branch (RGB) stars and 298,036 red clump (RC) stars, each with 6D phase-space coordinates, [Fe/H], $\alpha$/Fe], and age measurements. This study treats RGB and RC stars as a single population. We have also tested the MGAR trends separately for RGB and RC stars and found that the qualitative trends remain unchanged.
2.1 Quality Cuts and Orbital Parameters
The main source of uncertainty in the MGAR arises from stellar age measurements. To minimize the contamination from binary evolution or mass transfer, we remove stars with masses $0.7\,M_{\odot}$ (Wang et al., 2023a). Since the age estimates have a typical uncertainty floor of $\sim 20\%$ (Wang et al., 2023a), we exclude stars with age uncertainties exceeding 30%. We have verified that applying a stricter 25% cutoff yields similar results.
Additional quality criteria include signal-to-noise ratio $\mathrm{SNR}>20$, vertical distance $|Z|<10\,\mathrm{kpc}$, and Galactocentric radius $5\,\mathrm{kpc}<R<20\,\mathrm{kpc}$. We require flag_feh_apogee=0 and flag_afe_apogee=0 to exclude stars with unreliable [Fe/H] or $\alpha$/Fe] estimates outside the training grid. To avoid halo contamination, we retain only stars with $\mathrm{[Fe/H]}\geq-1.0$ dex and $v_{\phi}\geq 0$.
We further exclude stars with relative uncertainty >> 40% in any of the spatial or velocity parameters $X\in\{r,v_{r},v_{\phi},z,v_{z}\}$, i.e., we require
$\left|\dfrac{\Delta X}{X}\displaystyle\right|<40\%\text{ for all }X.$
These quality cuts result in a clean sample of 246,671 stars.
To mitigate the impact of orbital blurring, we follow V21 and compute gradients using the Galactocentric radius of the guiding center $R_{\mathrm{g}}$. $R_{\mathrm{g}}$ and the maximum vertical excursion $Z_{\mathrm{max}}$ are calculated using the Galactic potential model MWPotential2014, implemented in the galpy package (Bovy, 2015). We adopt the solar position $R_{0}=8.27\,\mathrm{kpc}$ and circular velocity $v_{0}=236\,\mathrm{km\,s^{-1}}$ (Schönrich & Aumer, 2017). Figure 1 shows the distributions and selection criteria (next Section) used for defining thin/thick disk subsamples.
Figure 1: Spatial and chemical distributions of the selected clean LAMOST DR8 sample. (a): Galactic meridional distribution. (b): Face-on view of the sample. (c): Number density in the $\alpha$/Fe]–[Fe/H] plane. The black line shows our chemical separation between the thin and thick disks ($\alpha$/Fe] =0.12=0.12 for [Fe/H] <0.3<-0.3, otherwise $\alpha$/Fe] $0.12-0.14(\mathrm{[Fe/H]}+0.3)$). The gray line indicates the V21 separation, which is not optimal for our sample. (d): Number density in the $Z_{\mathrm{max}}$$R_{\mathrm{g}}$ plane. The blue line at $Z_{\mathrm{max}}=0.25\,\mathrm{kpc}$ and red line at $Z_{\mathrm{max}}=2\,\mathrm{kpc}$ mark our geometric cuts for the thin and thick disk selections, respectively.
2.2 Thin/Thick Disk Subsamples
The MW’s disk can be separated into two components, either geometrically (via vertical excursion) or chemically (via $\alpha$/Fe]–[Fe/H] bimodality). Geometrically, the MW’s stellar density profile can be decomposed into two exponential components (e.g., Gilmore & Reid, 1983). In the $\mathrm{\alpha/Fe}][$$\mathrm{[Fe/H]}$ plane, there are two distinct sequences (e.g., Hayden et al., 2015); such a chemical bimodality reflects differences in nucleosynthetic timescales (Nomoto et al., 2013). The $\alpha$-rich sequence is associated with rapid early star formation dominated by Type II supernovae, which produce more $\alpha$ elements (e.g., $^{\text{16}}$O, $^{\text{20}}$Ne, $^{\text{24}}$Mg, $^{\text{28}}$Si) relative to $^{\text{56}}$Fe. In contrast, Type Ia supernovae, with their longer delay times, enrich more Fe relative to $\alpha$ and contribute more to the $\alpha$-poor (solar-like) sequence (Binney & Tremaine, 2008).
As argued by V21, using chemical and geometric cuts together more effectively isolates the flattening MGAR trend than applying either cut alone. We combine both criteria to select “pure” thin and thick disk populations as follows:
  • Thin disk: $\alpha$/Fe] below the black line in Figure 1(c) and $Z_{\mathrm{max}}<0.25\,\mathrm{kpc}$
  • Thick disk: $\alpha$/Fe] above the black line in Figure 1(c) and $Z_{\mathrm{max}}>2\,\mathrm{kpc}$.
This yields 18,363 thin disk stars and 24,210 thick disk stars (7.4% and 9.8% of the total clean sample, respectively). Their distributions in the $\alpha$/Fe]–[Fe/H] and $Z_{\mathrm{max}}$$R_{\mathrm{g}}$ planes are shown in Section 3.
[…]
2.3 Calculation of the MGAR
We calculate the metallicity gradient $\mathrm{d[Fe/H]}/\mathrm{d}R_{\mathrm{g}}$ across nine equal-width age bins from 1 to 12.25 Gyr. Following V21, we apply the weighted ordinary least squares (OLS) regression from SciPy (Virtanen et al., 2020) using weights $w_{i}=1/\Delta\mathrm{[Fe/H]}^{2}_{i}$, where $\Delta\mathrm{[Fe/H]}$ is the absolute metallicity uncertainty. The uncertainty in the fitted gradient is derived from the covariance matrix.
To evaluate the effect of uncertainties in [Fe/H], $R_{\mathrm{g}}$, and age simultaneously, we perform an independent Bayesian fit in Appendix A using the PyMC package (PyMC-Devs, 2025). This test incorporates all three sources of uncertainty and confirms that the MGAR trends in both disks are robust.
3 Results
Figure 3 presents the MGAR in the thin and thick disks, compared to the results of V21 (black points). Gradient fits for each mono-age population are shown in detail in Figure 4. Larger uncertainties—particularly for the oldest thin disk stars and the youngest thick disk populations—are primarily due to limited sample sizes (see the age histograms in Figure 2).
Figure 3: Metallicity gradient–age relation (MGAR) in the thin (top) and thick (bottom) disks, compared to the results of V21 (black dots). Error bars are derived from the covariance matrix of the weighted OLS fits. Upper panel: MGAR in the thin disk. The blue line shows our fitted trend, with a flattening rate of $0.0016\,\mathrm{dex\,kpc^{-1}\,Gyr^{-1}}$; the shaded region denotes the 98% highest density interval (HDI). The black line and gray shade show the fit for V21. Check the caption of Figure A1 in the Appendix for prior specifications. Lower panel: MGAR in the thick disk. Our results yield a mean positive gradient of $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$.
We also examine how the MGAR responds to variations in selection criteria and orbital integration configurations in Appendices B and C. Similar to V21, Appendix B shows that the flattening MGAR in the thin disk primarily originates from low-$Z_{\mathrm{max}}$ stars, which may be due to the favored migration for stars on the vertically coldest orbits (Solway et al., 2012; Vera-Ciro et al., 2014). Appendix C further demonstrates that the choice of Galactic potential $\Phi$ and solar circular velocity $v_{0}$ can affect the numerical values of MGAR in the thin disk, but the overall flattening trend remains robust.
3.1 Metallicity Gradient in the Thin Disk
The metallicity gradient in the thin disk gradually flattens as stellar age increases and aligns broadly with V21. The gradient of the youngest stars reaches $0.07\,\mathrm{dex\,kpc^{-1}}$ in agreement with V21 and other studies (Chen & Zhao, 2020; Zhang et al., 2021; Wang et al., 2023b; Ratcliffe et al., 2023).
A comprehensive comparison of thin disk MGAR measurements using various stellar tracers is given in Wang et al. (2023b). Their Figure 5 highlights a distinctively V-shaped MGAR trend with a dip around $5\sim 7\,\mathrm{Gyr}$ in studies based on main-sequence turn-off (MSTO) stars (e.g., Xiang et al., 2015; Wang et al., 2019), which is absent in our result using RGB and RC stars. V-shapes with a dip around 2 Gyr have been reported using red giant stars from CoRoGEE (Anders et al., 2017) and APOGEE (Anders et al., 2023). The existence and characteristics of this V-shape remain unclear and warrant further investigation with larger samples.
3.2 Metallicity Gradient in the Thick Disk
The gradients in the thick disk are generally positive across all ages, with a mean value of $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$. Interestingly, the gradient increases steadily with decreasing stellar age, reaching a peak of $\sim 0.019\,\mathrm{dex\,kpc^{-1}}$ at $\sim 5.4\,\mathrm{Gyr}$, but flattens at younger ages. This flattening is likely driven by a population of young, metal-poor thick disk stars. As shown in the fourth row of Figure 4, these stars—located at $R_{\mathrm{g}}>10\,\mathrm{kpc}$—contribute to suppressed gradients in the young thick disk panels. Our test confirms that thick disk stars at $R_{\mathrm{g}}>10\,\mathrm{kpc}$ exhibit a negative gradient ($\sim-0.02\,\mathrm{dex\,kpc^{-1}}$). Removing these stars reduces the flattening at young ages and raises the gradient values to $\sim 0.02\,\mathrm{dex\,kpc^{-1}}$, but the overall MGAR trend remains. Overall, our results are consistent with V21, reaffirming the existence of a slightly positive MGAR in the thick disk.
[…]
4 Discussion
Figure 5: Comparison between our observed MGAR in the thin disk and literature results (solid lines for MGARs, dashed lines for stellar birth/ISM gradient evolutions).Upper panel: MGAR in the thin disk compared with reconstructed birth gradients from Ratcliffe et al. (2023) and Lu et al. (2024). The divergence between the MGAR and birth gradient at old ages may reflect the cumulative effect of churning. The gradient of the youngest stars aligns with the birth gradients ($\sim-0.07\,\mathrm{dex\,kpc^{-1}}$) since little migration could occur. Lower panel: MGAR compared with two results predicting steepening gradients: the gas-phase evolution from Sharda et al. (2021b) and the birth gradient from FIRE2 simulations (Graf et al., 2024), we also plot the MGAR from the FIRE2 simulations. See the text for further discussion.
4.1 Origin of the Flattening Gradient in the Thin Disk
As discussed in Section 1, radial migration—particularly churning—is a likely mechanism for redistributing stars after their formation, thereby contributing to the observed flattening of the MGAR (see Figure 3, upper panel). However, it remains uncertain whether the flattening primarily reflects stellar migration or intrinsic evolution in the ISM gradient over time.
Several studies have explored this question. Within the inside-out formation framework, Minchev et al. (2018) reconstructed stellar birth radii and found a gradual flattening of the birth metallicity gradient over time. Similar flattening trends have been observed in external galaxies, where more massive systems tend to show shallower gradients. The analytic model of Sharda et al. (2021a) captures this mass dependence, predicting that gradients measured in $\mathrm{dex\,kpc^{-1}}$ flatten with increasing stellar mass, while gradients in $\mathrm{dex\,r_{e}^{-1}}$ follow a U-shape ($r_{e}$ is the effective radius). Interestingly, applying this model to MW–like galaxies yields opposite gradient evolutions depending on assumptions: a steepening ISM gradient over time, or a flattening gradient if the gas velocity dispersion is scaled by $(1+z)^{-1}$, to account for possible overestimates in star formation rates (Sharda et al., 2021b). These discrepancies highlight the need for more robust theoretical treatments.
Recent FIRE2 simulations challenged the migration-driven scenario. Graf et al. (2025) found that the present-day MGAR closely follows the evolution of the stellar birth gradient (or ISM gradient then), suggesting minimal impact from radial migration. As shown in the simulations of Graf et al. (2024), gas-phase metals are efficiently mixed by turbulence and diffusion before star formation, suppressing spatial metallicity differences. As a result, prominent metallicity gradients appear only after the disk settles and turbulence decreases. This interpretation emphasizes gas-phase processes over stellar redistribution.
Figure 5 contrasts our observed MGAR with several of the previously mentioned studies (solid lines for MGARs, dashed lines for stellar birth/ISM gradient evolutions). In the upper panel, we compare our thin disk MGAR with the reconstructed birth gradients from Ratcliffe et al. (2023) and Lu et al. (2024). The MGAR steepens and the birth gradients flatten as stellar age decreases, and they converge near $0.07\,\mathrm{dex\,kpc^{-1}}$ for the youngest stars, where migration is negligible. Lu et al. (2024) proposed that the sharp steepening of birth gradient between 12 and 10 Gyr in Ratcliffe et al. (2023) could be attributed to the Gaia–Enceladus/Sausage merger (Helmi et al., 2018; Belokurov et al., 2018).
In the lower panel of Figure 5, we compare our MGAR in the thin disk to the gas-phase gradient evolution predicted by Sharda et al. (2021b) and to both the birth gradient and MGAR from FIRE2 simulations (Graf et al., 2024). Our observed MGAR agrees broadly with Sharda et al. (2021b) except for the upturn at younger ages ($\lesssim 3$ Gyr), which is likely due to enhanced dilution from late gas accretion. Within FIRE2, the similarity between the MGAR (green solid line) and the birth gradient (green dashed line) suggests minimal churning. Similar to other studies, our MGAR appears significantly steeper than that of FIRE2, possibly due to uncertainties in Galactic distance measurements or an earlier disk settling in the MW compared to FIRE2 galaxies or nearby MW analogs (Graf et al., 2024).
4.2 Origin of the Positive Gradient in the Thick Disk
[…]
Given the relatively small positive gradient we observe ($\sim 0.013\,\mathrm{dex\,kpc^{-1}}$), another possibility is that the thick disk initially had no radial gradient. This would align with the clumpy formation scenario of the thick disk, where giant clumps heat the galactic disk and violent perturbations erase chemodynamical relations during the formation of the thick disk (e.g., Inoue & Saitoh, 2014).
The flattening gradient with increasing age for thick disk stars older than $\sim 5.4\,\mathrm{Gyr}$ might reflect the cumulative effects of radial migration within the thick disk (Solway et al., 2012; Halle et al., 2018). Early turbulence and mixing could have dispersed stars from their birth radii, weakening any original gradient and leading to the flatter distributions seen today.
Future studies could apply parameterized models such as the one used in Frankel et al. (2020) to improve the robustness of MGAR measurements. Such parameterized models and the multi-ring chemical evolution model in Johnson et al. (2021) may also help distinguish between different birth gradient evolution scenarios shown in Figure 5.
5 Conclusions
Following the methodology of Vickers et al. (2021, hereafter V21), we select chemically and kinematically “pure” thin and thick disk stars from LAMOST DR8 and compute the radial metallicity gradient ($\mathrm{d[Fe/H]}/\mathrm{d}R_{\mathrm{g}}$) across mono-age populations to investigate the metallicity gradient–age relation (MGAR) in both disks.
For the thin disk, we confirm a monotonic flattening of the MGAR with increasing stellar age. The fitted flattening rate is $\sim 0.0016\,\mathrm{dex\,kpc^{-1}\,Gyr^{-1}}$, which is lower than the value reported by V21 ($0.003\,\mathrm{dex\,kpc^{-1}\,Gyr^{-1}}$). The youngest thin disk stars exhibit a gradient of $0.07\,\mathrm{dex\,kpc^{-1}}$, consistent with V21 and other studies. In contrast, the thick disk shows a generally positive MGAR with a mean value of $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$, also in agreement with V21. Notably, we observe a flattening of the gradient for young thick disk stars (1.0–4.75 Gyr), likely due to a population of young, metal-poor stars at large guiding radii.
These MGAR trends reflect distinct chemodynamical histories for the MW’s thin and thick disks. In particular, the thin disk MGAR offers a valuable diagnostic for distinguishing between different formation scenarios, such as the “inside-out” model with significant radial migration and the “upside-down” model with minimal migration. Future studies using larger stellar samples with more precise ages, alongside improved chemodynamical modeling, will be crucial for disentangling these processes and refining our understanding of the Galaxy’s formation history.
6 ACKNOWLEDGEMENTS
We thank the developers and maintainers of the following software libraries, which were used in this work: ArviZ (Kumar et al., 2019), Astropy (Astropy Collaboration et al., 2022), galpy (Bovy, 2015), matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), pandas (pandas development team, 2020; Wes McKinney, 2010), PyMC (PyMC-Devs, 2025), seaborn (Waskom, 2021), and scipy (Virtanen et al., 2020).
The research presented here is partially supported by the National Natural Science Foundation of China under grant Nos. 12025302, 11773052, 11761131016; by China Manned Space Program with grant no. CMS-CSST-2025-A11; and by the “111” Project of the Ministry of Education of China under grant No. B20019. This work has made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.
7 Data Availability
The LAMOST DR8 astrometric and spectroscopic data, as well as the RGB and RC age catalog, are publicly available at: http://www.lamost.org/dr8/v1.0/doc/vac.
Appendix A Fitting MGAR Incorporating All Uncertainties
Figure A1: MGAR results from a Bayesian model that incorporates uncertainties in $\mathrm{[Fe/H]}$, guiding radius $R_{\mathrm{g}}$, and stellar age $\tau$. Left panel: MGARs in the thin disk. Blue points show the fiducial OLS results (this work), black points show V21, and forest green points are from the Bayesian model. Shaded regions denote the 98% highest density intervals (HDI) for each fitted line. All fits adopt the same prior: slope $\sim\mathcal{N}(0,2^{2})$, intercept $\sim\mathcal{N}(0,2^{2})$. Solid line fitting model: $\text{gradient}\sim\mathcal{N}(\text{intercept}+\text{slope}\times\text{age bin center},\epsilon^{2})$, where $\epsilon\sim\mathrm{HalfCauchy}(\beta=5)$. Right panel: MGARs in the thick disk: OLS fitting (red), V21 (black), and the Bayesian model (gold).
We assess how uncertainties in [Fe/H], guiding radius $R_{\mathrm{g}}$, and stellar age $\tau$ affect the derived MGAR by implementing a Bayesian fitting procedure using PyMC (PyMC-Devs, 2025). The absolute [Fe/H] uncertainty ($\Delta\mathrm{[Fe/H]}$) is taken directly from the LAMOST DR8 catalog. The absolute uncertainty in $R_{\mathrm{g}}$ ($\Delta R_{\mathrm{g}}$) is estimated by propagating position and velocity errors through 100 Monte Carlo realizations. The absolute stellar age uncertainty ($\Delta\tau$) is from Wang et al. (2023a), which includes an intrinsic method error of $\sim$ 20% (i.e., $\delta\tau\gtrsim 20\%$).
We model the relation $\mathrm{[Fe/H]}\sim\nabla\,R_{\mathrm{g}}+b$, where $\nabla$ is the metallicity gradient and $b$ is the intercept, assuming a truncated normal likelihood $\widetilde{\mathcal{N}}$:
$\mathrm{[Fe/H]}_{\mathrm{obs}}\sim\widetilde{\mathcal{N}}(\mathrm{[Fe/H]}\mid\nabla\,R_{\mathrm{g}}+b,\Sigma^{2}),$ (A1)
with the observed $\mathrm{[Fe/H]}_{\mathrm{obs}}$ truncated to the range [1.0, 0.5][-1.0,\,0.5] dex.
  • Uncertainty in $R_{\mathrm{g}}$: We treat $R_{\mathrm{g}}$ as a latent variable drawn from a truncated normal distribution:
    $R_{\mathrm{g}}\sim\widetilde{\mathcal{N}}(R_{\mathrm{g}}\mid R_{\mathrm{g,obs}},\Delta R_{\mathrm{g}}^{2}),\quad R_{\mathrm{g}}\in[0,\,20]~\mathrm{kpc}.$ (A2)
  • Uncertainty in [Fe/H]: The total variance $\Sigma$ combines intrinsic scatter $\sigma$ and absolute [Fe/H] uncertainty:
    $\Sigma^{2}\propto\sigma^{2}+\Delta\mathrm{[Fe/H]}^{2}.$ (A3)
  • Uncertainty in age: To account for $\Delta\tau$, we apply a weight based on the probability that the true age lies within the assigned age bin. Assuming a truncated normal distribution for the true age within $0,\,13]\,\mathrm{Gyr}$, the effective variance $\Sigma$ in Equation A1 becomes:
    $\Sigma^{2}=\dfrac{\sigma^{2}+\Delta\mathrm{[Fe/H]}^{2}}{\left[\displaystyle\int_{\mathrm{bin}}\widetilde{\mathcal{N}}(\tau\mid\tau_{\mathrm{obs}},\Delta\tau^{2})\,\mathrm{d}\tau\right]^{2}}.$ (A4)
  • Priors: The model uses the following priors:
    $\nabla\sim\mathcal{N}(0,0.5^{2}),\quad b\sim\mathcal{N}(0,0.5^{2}),\quad\sigma\sim\mathrm{HalfNormal}(0,0.4^{2}).$
Posterior sampling is performed using the PyMC package. The uncertainty of the gradient $\nabla$ is the standard deviation of the posterior.
Figure A1 compares the MGAR obtained from this Bayesian model to our fiducial (OLS) results and V21 results. We find that our Bayesian model reduces fluctuations in the MGAR for both the thin and thick disks. For the thin disk, the Bayesian MGAR trend shows the lowest flattening rate. In the thick disk, the Bayesian MGAR shifts upward, though the age-dependent pattern remains similar. These results confirm that the qualitative MGAR trends are robust when all key uncertainties are included.
[…]
We examine how changes to the chemical ($\alpha$/Fe]–[Fe/H]) and geometric (i.e., $Z_{\mathrm{max}}$) selection criteria affect the MGAR.
In Figure B1, we adjust the chemical division line in the $\alpha$/Fe]–[Fe/H] plane by $\pm 0.1\,\mathrm{dex}$ without applying the $Z_{\mathrm{max}}$ cut. The top row shows the resulting MGARs under these altered chemical selections.
Figure B2 explores the effect of purely geometric selections, dividing the sample into four $Z_{\mathrm{max}}$ bins: (0–0.25), (0.25–0.8), (0.8–2), and (2–6) kpc, with no chemical cut applied. The first and second columns of Figure B2 show that the observed MGAR flattening in the thin disk mainly comes from stars with low $Z_{\mathrm{max}}$, which may be due to the preferred migration for stars on the vertically coldest orbits (Solway et al., 2012; Vera-Ciro et al., 2014).
Figure B1: Effect of shifting the chemical division in the $\alpha$/Fe]–[Fe/H] plane by $\pm 0.1$ dex, blue and red dots are MGAR results from the main text. The top row displays the MGAR results without $Z_{\mathrm{max}}$ cut. The second row shows the chemical-only selection.
Figure B2: MGAR results for subsamples selected solely by $Z_{\mathrm{max}}$, with no chemical cut applied. The four columns correspond to $Z_{\mathrm{max}}$ = (0–0.25), (0.25–0.8), (0.8–2), and (2–6) kpc. The flattening MGAR trend in the thin disk is mainly contributed by stars with low $Z_{\mathrm{max}}$.
Appendix C Sensitivity to Galactic Potential and Solar Parameters
The canonical results in the main text use $\Phi=$MWPotential2014 as the Galactic potential, a solar Galactocentric radius of $R_{0}=8.27\,\mathrm{kpc}$, and a circular velocity of $v_{0}=236\,\mathrm{km\,s^{-1}}$ when calculating $R_{\mathrm{g}}$. To test the robustness of our results, we vary each of these parameters while keeping the others fixed, as shown in Figure C1.
In the first row, we adopt three different Galactic potentials available in galpy: MWPotential2014, McMillan17, and Irrgang13I. In the second row, we vary the solar radius $R_{0}$ across 8.27, 7.0, and 9.5 kpc. In the third row, we test $v_{0}=$ 236, 260, and 200 km s-1. The resulting MGARs remain largely consistent across these configurations. We find that while the numerical values of the MGAR in the thin disk are sensitive to the adopted $\Phi$ and $v_{0}$, the overall MGAR remains robust across all panels in Figure C1.
Figure C1: Robustness test of the MGAR under different Galactic configurations. The three rows correspond to variations in: (1) Galactic potential models used in galpy (MWPotential2014, McMillan17, Irrgang13I); (2) solar Galactocentric radius $R_{0}$ (8.27, 7.0, and 9.5 kpc); and (3) solar circular velocity $v_{0}$ (236, 260, and 200 km s-1). The canonical setup in the main text adopts $\Phi=$MWPotential2014, $R_{0}=8.27\,\mathrm{kpc}$, and $v_{0}=236\,\mathrm{km\,s^{-1}}$. The thin disk MGAR is most sensitive to $\Phi$ and $v_{0}$.

### v4

@@
 Recent studies report that this gradient becomes flatter with increasing stellar age in the Milky Way’s (MW) thin disk, while the thick disk exhibits a mildly positive gradient across all ages.
-In this work, we revisit the metallicity gradient–age relation (MGAR) in both the thin and thick disks of the MW.
-We use spectroscopic data from LAMOST DR8 and stellar ages calibrated with asteroseismology.
+In this work, we revisit the metallicity gradient–age relation (MGAR) in both the thin and thick disks of the MW, using spectroscopic data from LAMOST DR8 and stellar ages calibrated with asteroseismology.
 Our results show a steadily flattening MGAR in the thin disk and confirm a positive gradient $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$ in the thick disk.
@@
 These findings suggest distinct chemodynamical evolution paths for the MW’s thin and thick disks and provide valuable constraints for future models of Galactic chemical evolution.
-Milky Way dynamics (1051) — Milky Way disk (1050) — Milky Way evolution (1052)
+keywords:
+Galaxy: abundances – Galaxy: kinematics and dynamics – Galaxy: disc – Galaxy: evolution
+pubyear: 2025pagerange: Revisiting the Radial Metallicity Gradient-Age Relation in the Milky Way’s Thin and Thick DisksC
 1 Introduction
 A cornerstone of galaxy evolution is the negative radial metallicity gradient, which describes the decrease in stellar metallicity with increasing galactocentric radius.
-Matteucci & Francois (1989) reproduced this trend using “inside-out” formation models, in which the inner disk forms stars more rapidly and enriches earlier than the outer regions.
-This scenario naturally results in a negative metallicity gradient for stars formed from the interstellar medium (ISM, Matteucci, 2021).
-Following this framework, Minchev et al.
-(2018) found that the ISM gradient flattens over time as the outer disk becomes progressively enriched and approaches the central metallicity.
+Matteucci and Francois (1989) reproduced this trend using “inside-out” formation models, in which the inner disk forms stars more rapidly and enriches earlier than the outer regions.
+This scenario naturally results in a negative metallicity gradient for stars formed from the interstellar medium (ISM).
+Nowadays, the radial metallicity gradient has become a crucial diagnostic in different classes of models of Galactic evolution.
+First, chemical evolution models incorporate physical processes such as radial gas flows, star formation histories, and stellar yields to get a realistic metallicity gradient and study the formation of the MW (e.g., Grisoni et al.
+2018; Sharda et al.
+2021b; see Matteucci 2021 for a review).
+Second, descriptive models treat the gradient as a parameter to fit observational data and constrain the structure of the Galactic disk (e.g., Schönrich and Binney, 2009; Sanders and Binney, 2015; Frankel et al., 2020; Sharma et al., 2021).
+Finally, hybrid models combine chemical evolution with numerical simulations, as in Minchev et al.
+(2013) and Johnson et al.
+(2021).
+Utilizing the hybrid framework of Minchev et al.
+(2013), Minchev et al.
+(2018) inferred stellar birth radius and concluded that the ISM gradient flattens over time as the outer disk becomes progressively enriched and approaches the central metallicity.
 In this context, we refer to the stellar birth gradient as the radial metallicity gradient imprinted on stars at their formation, which inherits from the corresponding ISM gradient.
@@
 Using effective temperature ($T_{\mathrm{eff}}$) of main-sequence stars as a proxy for age, Yu et al.
-(2012) found that the gradient steepens with increasing $T_{\mathrm{eff}}$ (typically associated with decreasing age), but shows no trend for stars with $T_{\mathrm{eff}}<6,000\,\mathrm{K}$—those with main-sequence lifetimes $\gtrsim 10$ Gyr.
+(2012) found that the gradient steepens with increasing $T_{\mathrm{eff}}$ (typically associated with decreasing age), but shows no trend for stars with $T_{\mathrm{eff}}<6,000\,\mathrm{K}$ (i.e., those with main-sequence lifetimes $\gtrsim 10$ Gyr).
 These early findings hinted at a flattening gradient for increasingly older stars in the MW’s thin disk.
-Large-scale spectroscopic surveys have greatly facilitated precise age estimates, enabling direct MGAR measurements (e.g., Vickers et al., 2021; Wang et al., 2023b; Ratcliffe et al., 2023).
+Large-scale spectroscopic surveys have greatly facilitated precise age estimates, enabling direct MGAR measurements (e.g., Vickers et al., 2021; Wang et al., 2023b; Willett et al., 2023; Ratcliffe et al., 2023).
 Vickers et al.
 (2021, hereafter V21) used LAMOST DR5 and reported a flattening gradient in the thin disk, starting at $0.075\,\mathrm{dex\,kpc^{-1}}$ for the youngest stars and flattening with a rate of $0.003\,\mathrm{dex\,kpc^{-1}\,Gyr^{-1}}$.
-They also found a slightly positive gradient of $\sim 0.02\,\mathrm{dex\,kpc^{-1}}$ for all ages in the thick disk, consistent with constantly positive gradients of old, cool main-sequence stars ($T_{\mathrm{eff}}<6,000\,\mathrm{K}$) in Yu et al.
+Willett et al.
+(2023) used asteroseismic age and adopted a hierarchical Bayesian model for the gradient fitting.
+They found a similar flattening trend as V21 in the thin disk.
+In the thick disk, V21 found a slightly positive gradient of $\sim 0.02\,\mathrm{dex\,kpc^{-1}}$ across all age bins, consistent with constantly positive gradients of old, cool main-sequence stars ($T_{\mathrm{eff}}<6,000\,\mathrm{K}$) in Yu et al.
 (2012), which are likely from the thick disk.
-Radial migration has been proposed as a major process that alters stellar positions and flattens metallicity gradients (Sellwood & Binney, 2002).
+Radial migration has been proposed as a major process that alters stellar positions and flattens metallicity gradients (Sellwood and Binney, 2002).
 This mechanism, sometimes dubbed “churning”, causes significant radial mixing; non-axisymmetric perturbations such as transient spiral arms can induce large-scale stellar migration by exchanging stars across the corotation radius.
@@
 Churning flattens the MGAR as older stars have had more time to migrate, resulting in a flatter radial metallicity distribution.
-Recent simulations support the impact of radial migration.
-Lu et al.
-(2024) and Ratcliffe et al.
-(2023) analyzed the correlation between [Fe/H] spreads and stellar birth gradients in mono-age simulated populations, and recovered a time-dependent flattening of birth gradients.
+The metallicity gradients at stellar birth provided by the recent method [^1] https://github.com/BridgetRatcliffe/Rbirth for estimating the stellar birth radius ($R_{\mathrm{birth}}$) further supports the significance of radial migration.
+Building on the empirical framework of Minchev et al.
+(2018), a more systematic approach to infer $R_{\mathrm{birth}}$ from stellar ages and metallicities has been developed in the follow-up studies (Lu et al., 2024; Ratcliffe et al., 2023, 2025b, 2025a).
+In these works, the evolution of the stellar birth metallicity gradient is approximated as
+$\nabla\mathrm{[Fe/H]}(\tau)\approx\frac{\Delta\mathrm{[Fe/H]}(\tau)}{\Delta R(\tau)}\propto\frac{\mathrm{Range[Fe/H](age)}}{\Delta R(\tau)},$ (1)
+where $\tau$ is the lookback time and equals stellar age.
+A linear anti-correlation between $\nabla\mathrm{[Fe/H]}(\tau)$ and $\mathrm{Range[Fe/H](age)}$—validated with NIHAO-UHD (Buck et al., 2020) and HESTIA (Libeskind et al., 2020) simulations by Lu et al.
+(2024)—then provides an estimate of the birth gradient $\nabla\mathrm{[Fe/H]}(\tau)$ (with later refinements correcting $\Delta R(\tau)$ and accounting for selection effects; see Ratcliffe et al.
+2025b, a).
+Finally, the birth radius is obtained as
+$R_{\mathrm{birth}}(\mathrm{age},\mathrm{[Fe/H]})=\frac{\mathrm{[Fe/H]}-\mathrm{[Fe/H]}(R=0,\tau)}{\nabla\mathrm{[Fe/H]}(\tau)},$ (2)
+given an estimate of $\mathrm{[Fe/H]}(R=0,\tau)$.
+As a by-product of this method, the reconstructed stellar birth gradients $\nabla\mathrm{[Fe/H]}(\tau)$ become progressively flatter with decreasing lookback time/age.
 Anders et al.
-(2023) further compared observed MGAR from APOGEE DR17 with modeled ISM gradients evolution from Ratcliffe et al.
-(2023) to demonstrate the role of migration.
-However, the migration interpretation has been challenged by recent FIRE2 simulations (Graf et al., 2024, 2025).
-They found that present-day MGARs closely follow the stellar birth gradient with minimal influence from churning in these simulations.
-As gas-phase metal gets mixed before star formation, the role of churning is significantly weakened.
+(2023) further compared APOGEE MGAR measurements with the flattening birth-gradient evolution of Ratcliffe et al.
+(2023), demonstrating how radial migration could explain the difference between birth gradient evolution and observed MGAR trend.
+On the simulation side, models with smooth/bursty stellar feedback predict qualitatively different birth-gradient evolutions with redshift.
+As summarized in Garcia et al.
+(2025b), simulations with smooth feedback (such as EAGLE, Illustris, IllustrisTNG, and SIMBA) employ subgrid treatments of the ISM that produce relatively smooth metal redistribution and predict strong negative gradients flattening with time.
+In contrast, bursty-feedback simulations (e.g., FIRE2, SPICE Bursty, and Thesan Zoom) yield relatively shallower birth gradients at all epochs and better match current observations at high redshifts (Garcia et al., 2025a).
+In particular, recent FIRE2 simulations strongly challenge the migration-based interpretation of the MGAR (Graf et al., 2024, 2025).
+In these models, the present-day MGAR closely tracks the stellar birth gradients, implying minimal stellar radial migration in the simulations.
+And as an alternative to stellar migration, gas mixing (diffusion, advection, and turbulence) efficiently mixes gas-phase metals before forming new stars.
 Moreover, Graf et al.
-(2024) argued that vertical evolution—where turbulent gas settles over time, known as the “upside-down” scenario—may play a more dominant role than inside-out radial growth in shaping Galactic chemo-dynamical history.
-This suggests the need to consider vertical processes and possible differences in the evolutionary pathways of the thin and thick disks, as also done by V21.
+(2024) argued that vertical evolution, where turbulent gas gradually settles over time (“upside-down” disk formation), may play a more important role than inside-out radial growth in shaping the Galaxy’s chemo-dynamical history.
+This highlights the need to account for vertical processes and potentially distinct evolutionary pathways of the thin and thick disks, as also emphasized by V21.
 To constrain the MGAR in the MW’s thin and thick disks, we follow an approach similar to V21.
 We isolate “pure” thin and thick disk populations using both geometric and chemical criteria, and compute MGAR separately for each population.
-Stellar kinematics and abundances are taken from LAMOST DR8 (Wang et al., 2022a), and stellar ages are calibrated from asteroseismic data provided by the Kepler mission (Wang et al., 2023a), offering an alternative to traditional isochrone fitting.
+Stellar kinematics and abundances are taken from LAMOST DR8 (Wang et al., 2022a), and stellar ages are calibrated from asteroseismic data provided by the Kepler mission (Gilliland et al., 2010; Borucki et al., 2010; Wang et al., 2023a), offering an alternative to traditional isochrone fitting.
 This paper is organized as follows.
@@
 2 Data and Method
-The Large Area Multi-Object Spectroscopic Telescope (LAMOST), located at Xinglong Station of the National Astronomical Observatory of China, is a 4-meter Schmidt telescope with a 5° field of view.
+The Large Area Multi-Object Spectroscopic Telescope (LAMOST, Deng et al., 2012; Zhao et al., 2012; Liu et al., 2014), located at Xinglong Station of the National Astronomical Observatory of China, is a 4-meter Schmidt telescope with a 5° field of view.
 It is one of the most extensive stellar spectroscopic surveys and has released 13 data sets to date.
 In this work, we use LAMOST Data Release 8 (DR8; Wang et al., 2022a) combined with stellar ages from Wang et al.
-(2023a), which trained a neural network using asteroseismic data from the LAMOST-Kepler survey.
-Our cross-matched sample contains 986,405 unique stars: 688,369 red giant branch (RGB) stars and 298,036 red clump (RC) stars, each with 6D phase-space coordinates, [Fe/H], $\alpha$/Fe], and age measurements.
+(2023a), which trained a neural network using asteroseismic data from the LAMOST-Kepler common stars.
+Our cross-matched sample comprises 986,405 unique stars, including 688,369 red giant branch (RGB) stars and 298,036 red clump (RC) stars, each with 6D phase-space coordinates, [Fe/H], $\alpha$/Fe], and age measurements.
 This study treats RGB and RC stars as a single population.
@@
 $R_{\mathrm{g}}$ and the maximum vertical excursion $Z_{\mathrm{max}}$ are calculated using the Galactic potential model MWPotential2014, implemented in the galpy package (Bovy, 2015).
-We adopt the solar position $R_{0}=8.27\,\mathrm{kpc}$ and circular velocity $v_{0}=236\,\mathrm{km\,s^{-1}}$ (Schönrich & Aumer, 2017).
+We adopt the solar position $R_{0}=8.27\,\mathrm{kpc}$ and circular velocity $v_{0}=236\,\mathrm{km\,s^{-1}}$ (Schönrich and Aumer, 2017).
 Figure 1 shows the distributions and selection criteria (next Section) used for defining thin/thick disk subsamples.
@@
 The blue line at $Z_{\mathrm{max}}=0.25\,\mathrm{kpc}$ and red line at $Z_{\mathrm{max}}=2\,\mathrm{kpc}$ mark our geometric cuts for the thin and thick disk selections, respectively.
+[…]
 2.2 Thin/thick disk subsamples
 The MW’s disk can be separated into two components, either geometrically (via vertical excursion) or chemically (via $\alpha$/Fe]–[Fe/H] bimodality).
-Geometrically, the MW’s stellar density profile can be decomposed into two exponential components (e.g., Gilmore & Reid, 1983).
-In the $\mathrm{\alpha/Fe}][$$\mathrm{[Fe/H]}$ plane, there are two distinct sequences (e.g., Hayden et al., 2015); such a chemical bimodality reflects differences in nucleosynthetic timescales (Nomoto et al., 2013).
+Geometrically, the MW’s stellar density profile can be decomposed into two exponential components (e.g., Gilmore and Reid, 1983).
+In the $\mathrm{\alpha/Fe}][$$\mathrm{[Fe/H]}$ plane, there are two distinct sequences (e.g., Hayden et al., 2015; Imig et al., 2023); such a chemical bimodality reflects differences in nucleosynthetic timescales (Nomoto et al., 2013).
 The $\alpha$-rich sequence is associated with rapid early star formation dominated by Type II supernovae, which produce more $\alpha$ elements (e.g., $^{\text{16}}$O, $^{\text{20}}$Ne, $^{\text{24}}$Mg, $^{\text{28}}$Si) relative to $^{\text{56}}$Fe.
-In contrast, Type Ia supernovae, with their longer delay times, enrich more Fe relative to $\alpha$ and contribute more to the $\alpha$-poor (solar-like) sequence (Binney & Tremaine, 2008).
+In contrast, Type Ia supernovae, with their longer delay times, enrich more Fe relative to $\alpha$ and contribute more to the $\alpha$-poor (solar-like) sequence (Binney and Tremaine, 2008).
 As argued by V21, using chemical and geometric cuts together more effectively isolates the flattening MGAR trend than applying either cut alone.
 We combine both criteria to select “pure” thin and thick disk populations as follows:
+  1. 1.
     Thin disk: $\alpha$/Fe] below the black line in Figure 1(c) and $Z_{\mathrm{max}}<0.25\,\mathrm{kpc}$
+  2. 2.
     Thick disk: $\alpha$/Fe] above the black line in Figure 1(c) and $Z_{\mathrm{max}}>2\,\mathrm{kpc}$.
@@
 […]
-[…]
+Figure 3: Metallicity gradient–age relation (MGAR) in the thin (top) and thick (bottom) disks, compared to the results of V21 (black dots).
+Error bars are derived from the covariance matrix of the weighted OLS fits.
+Upper panel: MGAR in the thin disk.
+The blue line shows our fitted trend, with a flattening rate of $0.0016\,\mathrm{dex\,kpc^{-1}\,Gyr^{-1}}$; the shaded region denotes the 98% highest density interval (HDI).
+The black line and gray shade show the fit for V21.
+Check the caption of Figure A1 in the Appendix for prior specifications.
+Lower panel: MGAR in the thick disk.
+Our results yield a mean positive gradient of $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$.
 2.3 Calculation of the MGAR
@@
 Gradient fits for each mono-age population are shown in detail in Figure 4.
-Larger uncertainties—particularly for the oldest thin disk stars and the youngest thick disk populations—are primarily due to limited sample sizes (see the age histograms in Figure 2).
-Figure 3: Metallicity gradient–age relation (MGAR) in the thin (top) and thick (bottom) disks, compared to the results of V21 (black dots).
-Error bars are derived from the covariance matrix of the weighted OLS fits.
-Upper panel: MGAR in the thin disk.
-The blue line shows our fitted trend, with a flattening rate of $0.0016\,\mathrm{dex\,kpc^{-1}\,Gyr^{-1}}$; the shaded region denotes the 98% highest density interval (HDI).
-The black line and gray shade show the fit for V21.
-Check the caption of Figure A1 in the Appendix for prior specifications.
-Lower panel: MGAR in the thick disk.
-Our results yield a mean positive gradient of $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$.
+Larger uncertainties for the oldest thin disk stars and the youngest thick disk populations are primarily due to limited sample sizes (see the age histograms in Figure 2).
 We also examine how the MGAR responds to variations in selection criteria and orbital integration configurations in Appendices B and C.
 Similar to V21, Appendix B shows that the flattening MGAR in the thin disk primarily originates from low-$Z_{\mathrm{max}}$ stars, which may be due to the favored migration for stars on the vertically coldest orbits (Solway et al., 2012; Vera-Ciro et al., 2014).
+For the intermediate $Z_{\mathrm{max}}$, we observe a slightly non-monotonic MGAR, similar to V21 and Willett et al.
+(2023).
 Appendix C further demonstrates that the choice of Galactic potential $\Phi$ and solar circular velocity $v_{0}$ can affect the numerical values of MGAR in the thin disk, but the overall flattening trend remains robust.
 3.1 Metallicity gradient in the thin disk
-The metallicity gradient in the thin disk gradually flattens as stellar age increases and aligns broadly with V21.
-The gradient of the youngest stars reaches $0.07\,\mathrm{dex\,kpc^{-1}}$ in agreement with V21 and other studies (Chen & Zhao, 2020; Zhang et al., 2021; Wang et al., 2023b; Ratcliffe et al., 2023).
+The metallicity gradient in the thin disk gradually flattens as stellar age increases and aligns broadly with V21 and Willett et al.
+(2023).
+The gradient of the youngest stars reaches $0.07\,\mathrm{dex\,kpc^{-1}}$ in agreement with V21 and other studies (Chen and Zhao, 2020; Zhang et al., 2021; Wang et al., 2023b; Willett et al., 2023; Ratcliffe et al., 2023; Sun et al., 2024).
 A comprehensive comparison of thin disk MGAR measurements using various stellar tracers is given in Wang et al.
@@
 V-shapes with a dip around 2 Gyr have been reported using red giant stars from CoRoGEE (Anders et al., 2017) and APOGEE (Anders et al., 2023).
-The existence and characteristics of this V-shape remain unclear and warrant further investigation with larger samples.
+It is worth noting that the V-shapes seem to appear more frequently when the gradient is calculated with respect to $R_{\mathrm{present}}$ (Xiang et al., 2015; Anders et al., 2017; Wang et al., 2019) rather than $R_{\mathrm{g}}$ (V21; Willett et al., 2023), though the V-shape of MGAR in Anders et al.
+(2023) is not sensitive to either radius.
+The existence and characteristics of these V-shapes remain unclear and warrant further investigation with larger samples.
+[…]
 3.2 Metallicity gradient in the thick disk
 The gradients in the thick disk are generally positive across all ages, with a mean value of $\sim 0.013\,\mathrm{dex\,kpc^{-1}}$.
-Interestingly, the gradient increases steadily with decreasing stellar age, reaching a peak of $\sim 0.019\,\mathrm{dex\,kpc^{-1}}$ at $\sim 5.4\,\mathrm{Gyr}$, but flattens at younger ages.
+Li and Zhao (2017) reported a higher value of $0.035\,\mathrm{dex\,kpc^{-1}}$ using 2,035 thick-disk giant stars from LAMOST DR3.
+Recently, Hu et al.
+(2023) measured the metallicity gradient with respect to the angular momentum $L_{z}$ in their chemically thick (high-$\alpha$ metal-poor) disk: $1.2\times 10^{-3}\,\mathrm{dex\,kpc^{-1}\,km^{-1}\,s}$, which corresponds to $\mathrm{d[M/H]}/\mathrm{d}R_{\mathrm{g}}=0.28\,\mathrm{dex\,kpc^{-1}}$ and is significantly larger than our findings here.
+Sun et al.
+(2024) found a positive gradient weaker than $\sim 0.01\,\mathrm{dex\,kpc^{-1}}$ in the high-$\alpha$ disk across different $|Z|$ bins.
+Interestingly, our gradient increases steadily with decreasing stellar age, reaching a peak of $\sim 0.019\,\mathrm{dex\,kpc^{-1}}$ at $\sim 5.4\,\mathrm{Gyr}$, but flattens at younger ages.
 This flattening is likely driven by a population of young, metal-poor thick disk stars.
-As shown in the fourth row of Figure 4, these stars—located at $R_{\mathrm{g}}>10\,\mathrm{kpc}$—contribute to suppressed gradients in the young thick disk panels.
+As shown in the fourth row of Figure 4, the stars located at $R_{\mathrm{g}}>10\,\mathrm{kpc}$ help to suppress gradients in the young thick disk panels.
 Our test confirms that thick disk stars at $R_{\mathrm{g}}>10\,\mathrm{kpc}$ exhibit a negative gradient ($\sim-0.02\,\mathrm{dex\,kpc^{-1}}$).
@@
 Overall, our results are consistent with V21, reaffirming the existence of a slightly positive MGAR in the thick disk.
-[…]
 4 Discussion
-Figure 5: Comparison between our observed MGAR in the thin disk and literature results (solid lines for MGARs, dashed lines for stellar birth/ISM gradient evolutions).Upper panel: MGAR in the thin disk compared with reconstructed birth gradients from Ratcliffe et al.
-(2023) and Lu et al.
-(2024).
+4.1 Origin of the flattening gradient in the thin disk
+As discussed in Section 1, radial migration (particularly churning) is a likely mechanism for redistributing stars after their formation, thereby contributing to the observed flattening of the MGAR (see Figure 3, upper panel).
+However, it remains uncertain whether the flattening primarily reflects stellar migration or intrinsic evolution in the ISM gradient over time.
+Figure 5: Comparison between our observed MGAR in the thin disk and literature results (solid lines for MGARs, dashed/dash-dotted lines for stellar birth/gas-phase gradient evolutions).
+Upper panel: MGAR in the thin disk compared with reconstructed birth gradients retrieved from Rbirth python package (calculated from age data from Anders et al.
+(2023) and the selection function corrected results from Khoperskov et al.
+(2025)).
 The divergence between the MGAR and birth gradient at old ages may reflect the cumulative effect of churning.
-The gradient of the youngest stars aligns with the birth gradients ($\sim-0.07\,\mathrm{dex\,kpc^{-1}}$) since little migration could occur.
-Lower panel: MGAR compared with two results predicting steepening gradients: the gas-phase evolution from Sharda et al.
-(2021b) and the birth gradient from FIRE2 simulations (Graf et al., 2024), we also plot the MGAR from the FIRE2 simulations.
+The gradient of the youngest stars aligns with the birth gradients ($\sim-0.065\,\mathrm{dex\,kpc^{-1}}$), as little migration is expected to occur.
+Lower panel: MGAR compared with steepening gradients: the gas-phase metallicity gradient evolution model from Sharda et al.
+(2021b, dash-dotted lines and the grey shade.
+$\phi_{y}$ is the yield reduction factor of the model: larger $\phi_{y}$ for stronger ISM enrichment; smaller $\phi_{y}$ for metallicity losses through galactic winds.) and the birth gradient from FIRE2 simulations (Graf et al., 2024), we also plot the MGAR from the FIRE2 simulations.
 See the text for further discussion.
-4.1 Origin of the Flattening Gradient in the Thin Disk
-As discussed in Section 1, radial migration—particularly churning—is a likely mechanism for redistributing stars after their formation, thereby contributing to the observed flattening of the MGAR (see Figure 3, upper panel).
-However, it remains uncertain whether the flattening primarily reflects stellar migration or intrinsic evolution in the ISM gradient over time.
 Several studies have explored this question.
-Within the inside-out formation framework, Minchev et al.
-(2018) reconstructed stellar birth radii and found a gradual flattening of the birth metallicity gradient over time.
+Recent method for estimating $R_{\mathrm{birth}}$ introduced in Section 1 found a gradually flattening of the birth metallicity gradients with time (Ratcliffe et al., 2025a).
 Similar flattening trends have been observed in external galaxies, where more massive systems tend to show shallower gradients.
 The analytic model of Sharda et al.
-(2021a) captures this mass dependence, predicting that gradients measured in $\mathrm{dex\,kpc^{-1}}$ flatten with increasing stellar mass, while gradients in $\mathrm{dex\,r_{e}^{-1}}$ follow a U-shape ($r_{e}$ is the effective radius).
-Interestingly, applying this model to MW–like galaxies yields opposite gradient evolutions depending on assumptions: a steepening ISM gradient over time, or a flattening gradient if the gas velocity dispersion is scaled by $(1+z)^{-1}$, to account for possible overestimates in star formation rates (Sharda et al., 2021b).
+(2021a) captures this mass dependence, predicting that gas-phase metallicity gradients measured in $\mathrm{dex\,kpc^{-1}}$ flatten with increasing stellar mass, while gradients in $\mathrm{dex\,r_{e}^{-1}}$ follow a U-shape ($r_{e}$ is the effective radius).
+Interestingly, applying this model to MW–like galaxies yields opposite gradient evolutions depending on assumptions: a steepening ISM gradient over time in their Figure 10, or a flattening gradient in their Figure B1 if the gas velocity dispersion is reduced by $(1+z)$, to account for possible overestimates in star formation rates (Sharda et al., 2021b).
 These discrepancies highlight the need for more robust theoretical treatments.
@@
 (2024), gas-phase metals are efficiently mixed by turbulence and diffusion before star formation, suppressing spatial metallicity differences.
-As a result, prominent metallicity gradients appear only after the disk settles and turbulence decreases.
+As a result, prominent metallicity gradients emerge only after the disk reaches a settled state and turbulence diminishes.
 This interpretation emphasizes gas-phase processes over stellar redistribution.
-Figure 5 contrasts our observed MGAR with several of the previously mentioned studies (solid lines for MGARs, dashed lines for stellar birth/ISM gradient evolutions).
-In the upper panel, we compare our thin disk MGAR with the reconstructed birth gradients from Ratcliffe et al.
-(2023) and Lu et al.
+Figure 5 contrasts our observed MGAR with several previously mentioned studies: solid lines for MGARs, dashed lines for stellar birth–gradient evolutions, and dot–dashed lines for gas–phase gradient evolutions.
+4.1.1 Comparison with flattening birth gradient
+In the upper panel of Figure 5, we compare our thin–disk MGAR with two reconstructed birth–gradient evolution trends retrieved from the Rbirth Python package, one based on the ages from Anders et al.
+(2023), another from the selection–function–corrected results of Khoperskov et al.
 (2025).
-The MGAR steepens and the birth gradients flatten as stellar age decreases, and they converge near $0.07\,\mathrm{dex\,kpc^{-1}}$ for the youngest stars, where migration is negligible.
-Lu et al.
-(2024) proposed that the sharp steepening of birth gradient between 12 and 10 Gyr in Ratcliffe et al.
-(2023) could be attributed to the Gaia–Enceladus/Sausage merger (Helmi et al., 2018; Belokurov et al., 2018).
-In the lower panel of Figure 5, we compare our MGAR in the thin disk to the gas-phase gradient evolution predicted by Sharda et al.
-(2021b) and to both the birth gradient and MGAR from FIRE2 simulations (Graf et al., 2024).
-Our observed MGAR agrees broadly with Sharda et al.
-(2021b) except for the upturn at younger ages ($\lesssim 3$ Gyr), which is likely due to enhanced dilution from late gas accretion.
-Within FIRE2, the similarity between the MGAR (green solid line) and the birth gradient (green dashed line) suggests minimal churning.
-Similar to other studies, our MGAR appears significantly steeper than that of FIRE2, possibly due to uncertainties in Galactic distance measurements or an earlier disk settling in the MW compared to FIRE2 galaxies or nearby MW analogs (Graf et al., 2024).
+The upper panel reveals a growing divergence toward older ages (i.e., larger lookback time): the reconstructed birth gradients steepen substantially, whereas our observed MGAR becomes progressively flatter.
+As a result, the difference between the two reaches $\geq 0.08\,\mathrm{dex\,kpc^{-1}}$ by $\sim 10\,\mathrm{Gyr}$.
+This widening offset at old ages is most naturally interpreted as the cumulative effect of radial migration: stars born with a steep radial metallicity gradient can be redistributed over large radial ranges over several Gyrs, thereby flattening the present–day MGAR relative to the (much steeper) reconstructed birth gradients.
+Although the reconstructed curves also carry uncertainties and may be affected by systematics, the increased discrepancy toward older ages implies that migration is an important shaping mechanism for the oldest thin disk populations.
+We further note that Lu et al.
+(2024) associated the sharp steepening in the reconstructed gradients around 10–12 Gyr with the Gaia–Enceladus/Sausage merger (Helmi et al., 2018; Belokurov et al., 2018).
+This event adds an additional complication when comparing reconstructed birth gradients with the present–day MGAR at large lookback times, but does not alter the qualitative conclusion that migration helps explain the observed divergence.
+4.1.2 Comparison with steepening birth/gas gradient
+In the lower panel of Figure 5, we compare the MGAR of our thin disk with both the MGAR and the birth gradients from FIRE2 simulations (Graf et al., 2024), and with the analytical evolution of the gas–phase metallicity gradient from Sharda et al.
+(2021b).
+In the analytical model, the parameter $\phi_{y}$ ($0\leq\phi_{y}\leq 1$) represents the yield reduction factor: larger $\phi_{y}$ values imply that most supernova-produced metals remain in the ISM, while smaller $\phi_{y}$ values correspond to stronger metallicity losses through galactic winds.
+Our observed MGAR broadly follows the $\phi_{y}=1.0$ model for ages $\gtrsim 3$ Gyr, except for an upturn at younger ages ($\lesssim 3$ Gyr), which is likely caused by dilution from late gas accretion.
+This agreement at intermediate and old ages suggests that, in this regime, the observed MGAR largely tracks the long–term evolution of the ISM metallicity gradient, and that radial migration plays only a secondary role.
+However, because the Sharda et al.
+(2021b) model assumes cylindrical symmetry and omits the bar and spiral structures, which are key drivers of radial migration in the Milky Way (Sellwood and Binney, 2002; Roskar, 2010; Zhang et al., 2025; Marques et al., 2025), hence its predicted gas-phase gradients are likely steeper than in a fully realistic disk.
+Thus, the apparent agreement with the $\phi_{y}=1.0$ curve should be interpreted with caution.
+For the FIRE2 simulations, both the MGAR and the birth gradients (green solid/dashed lines) lie between the $\phi_{y}=0.1$ and $\phi_{y}=0.3$ curves, substantially flatter than our observed MGAR, which is much closer to the $\phi_{y}=1.0$ curve.
+Thus, the vertical offset is plausibly explained by the stronger metal ejection in FIRE2, which consequently produces flatter ISM/birth gradients.
+Other possible contributors to the offset include uncertainties in distance measurements or the possibility that the real Milky Way experienced earlier disk settling than FIRE2 galaxies or nearby analogs (Graf et al., 2024).
+Although the overall trends of MGAR and birth gradients in FIRE2 are similar (i.e., the agreement between the green solid line and green dashed line), which indicates limited radial migration within the simulations, the weak and short bars in FIRE2 (Ansar et al., 2025) suggest that the efficiency of churning may be underestimated relative to the real Milky Way.
+Therefore, the FIRE2 agreement between MGAR and the birth gradients reflects the underlying ISM evolution in the simulations rather than providing a direct constraint on the Milky Way’s migration efficiency.
 4.2 Origin of the positive gradient in the thick disk
-[…]
+The thick disk likely formed in the early evolutionary stages of the MW, when the disk was gas-rich and dynamically turbulent.
+Under such conditions, it is plausible that a weakly positive radial metallicity gradient could emerge.
+Simulations from the FIRE2 suite suggest that strong stellar feedback, especially from early starbursts, can drive metal-enriched outflows into the outer disk, leading to such positive gradients (Sun et al., 2025).
+The accretion of pristine gas into the MW center during early epochs may also contribute to forming a positive metallicity gradient (Cresci et al., 2010; Curir et al., 2012).
+Recent JWST observations show positive metallicity gradient ($0.165\,\mathrm{dex\,kpc^{-1}}$) at z $\sim 3$ (Wang et al., 2022b).
+For even higher redshifts, Venturi et al.
+(2024) measured slightly positive gradients ($0.03\sim 0.14\,\mathrm{dex\,kpc^{-1}}$) in $6<z<8$, while Li et al.
+(2025) reported flattening negative gradients ($0.53\,\mathrm{dex\,kpc^{-1}}$ to $0.34\,\mathrm{dex\,kpc^{-1}}$) from z $\approx 7$ to z $\approx 6$ (their Figure 2) and suggested that the slightly positive gradients in Venturi et al.
+(2024) are resulted from merging systems.
 Given the relatively small positive gradient we observe ($\sim 0.013\,\mathrm{dex\,kpc^{-1}}$), another possibility is that the thick disk initially had no radial gradient.
-This would align with the clumpy formation scenario of the thick disk, where giant clumps heat the galactic disk and violent perturbations erase chemodynamical relations during the formation of the thick disk (e.g., Inoue & Saitoh, 2014).
+This would align with the clumpy formation scenario of the thick disk, where giant clumps heat the galactic disk and violent perturbations erase chemodynamical relations during the formation of the thick disk (e.g., Inoue and Saitoh, 2014).
 The flattening gradient with increasing age for thick disk stars older than $\sim 5.4\,\mathrm{Gyr}$ might reflect the cumulative effects of radial migration within the thick disk (Solway et al., 2012; Halle et al., 2018).
 Early turbulence and mixing could have dispersed stars from their birth radii, weakening any original gradient and leading to the flatter distributions seen today.
-Future studies could apply parameterized models such as the one used in Frankel et al.
-(2020) to improve the robustness of MGAR measurements.
-Such parameterized models and the multi-ring chemical evolution model in Johnson et al.
-(2021) may also help distinguish between different birth gradient evolution scenarios shown in Figure 5.
+Future studies could aim to correct for the selection function and apply parameterized models to improve the robustness of MGAR measurements.
+Such parameterized models and chemical evolution models mentioned in Section 1 may also help to distinguish between different birth gradient evolution scenarios shown in Figure 5.
 5 Conclusions
@@
 Future studies using larger stellar samples with more precise ages, alongside improved chemodynamical modeling, will be crucial for disentangling these processes and refining our understanding of the Galaxy’s formation history.
-6 ACKNOWLEDGEMENTS
-We thank the developers and maintainers of the following software libraries, which were used in this work: ArviZ (Kumar et al., 2019), Astropy (Astropy Collaboration et al., 2022), galpy (Bovy, 2015), matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), pandas (pandas development team, 2020; Wes McKinney, 2010), PyMC (PyMC-Devs, 2025), seaborn (Waskom, 2021), and scipy (Virtanen et al., 2020).
+Acknowledgements
+We thank the anonymous referee for helpful and constructive suggestions that strengthened the quality of this work.
+We thank the developers and maintainers of the following software libraries, which were used in this work: ArviZ (Kumar et al., 2019), Astropy (Astropy Collaboration et al., 2022), galpy (Bovy, 2015), matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), pandas (pandas development team, 2020; McKinney, 2010), PyMC (PyMC-Devs, 2025), seaborn (Waskom, 2021), and scipy (Virtanen et al., 2020).
 The research presented here is partially supported by the National Natural Science Foundation of China under grant Nos.
-12025302, 11773052, 11761131016; by China Manned Space Program with grant no.
+12533004, 12025302, 11773052; by China Manned Space Program with grant no.
 CMS-CSST-2025-A11; and by the “111” Project of the Ministry of Education of China under grant No.
@@
 This work has made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.
-7 Data Availability
+Data Availability
 The LAMOST DR8 astrometric and spectroscopic data, as well as the RGB and RC age catalog, are publicly available at: http://www.lamost.org/dr8/v1.0/doc/vac.
@@
 We model the relation $\mathrm{[Fe/H]}\sim\nabla\,R_{\mathrm{g}}+b$, where $\nabla$ is the metallicity gradient and $b$ is the intercept, assuming a truncated normal likelihood $\widetilde{\mathcal{N}}$:
-$\mathrm{[Fe/H]}_{\mathrm{obs}}\sim\widetilde{\mathcal{N}}(\mathrm{[Fe/H]}\mid\nabla\,R_{\mathrm{g}}+b,\Sigma^{2}),$ (A1)
+$\mathrm{[Fe/H]}_{\mathrm{obs}}\sim\widetilde{\mathcal{N}}(\mathrm{[Fe/H]}\mid\nabla\,R_{\mathrm{g}}+b,\Sigma^{2}),$ (3)
 with the observed $\mathrm{[Fe/H]}_{\mathrm{obs}}$ truncated to the range [1.0, 0.5][-1.0,\,0.5] dex.
+  1. 1.
     Uncertainty in $R_{\mathrm{g}}$: We treat $R_{\mathrm{g}}$ as a latent variable drawn from a truncated normal distribution:
-    $R_{\mathrm{g}}\sim\widetilde{\mathcal{N}}(R_{\mathrm{g}}\mid R_{\mathrm{g,obs}},\Delta R_{\mathrm{g}}^{2}),\quad R_{\mathrm{g}}\in[0,\,20]~\mathrm{kpc}.$ (A2)
+    $R_{\mathrm{g}}\sim\widetilde{\mathcal{N}}(R_{\mathrm{g}}\mid R_{\mathrm{g,obs}},\Delta R_{\mathrm{g}}^{2}),\quad R_{\mathrm{g}}\in[0,\,20]~\mathrm{kpc}.$ (4)
+  2. 2.
     Uncertainty in [Fe/H]: The total variance $\Sigma$ combines intrinsic scatter $\sigma$ and absolute [Fe/H] uncertainty:
-    $\Sigma^{2}\propto\sigma^{2}+\Delta\mathrm{[Fe/H]}^{2}.$ (A3)
+    $\Sigma^{2}\propto\sigma^{2}+\Delta\mathrm{[Fe/H]}^{2}.$ (5)
+  3. 3.
     Uncertainty in age: To account for $\Delta\tau$, we apply a weight based on the probability that the true age lies within the assigned age bin.
-Assuming a truncated normal distribution for the true age within $0,\,13]\,\mathrm{Gyr}$, the effective variance $\Sigma$ in Equation A1 becomes:
-    $\Sigma^{2}=\dfrac{\sigma^{2}+\Delta\mathrm{[Fe/H]}^{2}}{\left[\displaystyle\int_{\mathrm{bin}}\widetilde{\mathcal{N}}(\tau\mid\tau_{\mathrm{obs}},\Delta\tau^{2})\,\mathrm{d}\tau\right]^{2}}.$ (A4)
+Assuming a truncated normal distribution for the true age within $0,\,13]\,\mathrm{Gyr}$, the effective variance $\Sigma$ in Equation 3 becomes:
+    $\Sigma^{2}=\dfrac{\sigma^{2}+\Delta\mathrm{[Fe/H]}^{2}}{\left[\displaystyle\int_{\mathrm{bin}}\widetilde{\mathcal{N}}(\tau\mid\tau_{\mathrm{obs}},\Delta\tau^{2})\,\mathrm{d}\tau\right]^{2}}.$ (6)
+  4. 4.
     Priors: The model uses the following priors:
-    $\nabla\sim\mathcal{N}(0,0.5^{2}),\quad b\sim\mathcal{N}(0,0.5^{2}),\quad\sigma\sim\mathrm{HalfNormal}(0,0.4^{2}).$
+    $\nabla\sim\mathcal{N}(0,0.5^{2}),\,b\sim\mathcal{N}(0,0.5^{2}),\,\sigma\sim\mathrm{HalfNormal}(0,0.4^{2}).$ (7)
 Posterior sampling is performed using the PyMC package. The uncertainty of the gradient $\nabla$ is the standard deviation of the posterior.
@@
 The first and second columns of Figure B2 show that the observed MGAR flattening in the thin disk mainly comes from stars with low $Z_{\mathrm{max}}$, which may be due to the preferred migration for stars on the vertically coldest orbits (Solway et al., 2012; Vera-Ciro et al., 2014).
+The slightly non-monotonic MGAR of stars with intermediate $Z_{\mathrm{max}}$ is also shown in V21 and Willett et al.
+(2023).
 Figure B1: Effect of shifting the chemical division in the $\alpha$/Fe]–[Fe/H] plane by $\pm 0.1$ dex, blue and red dots are MGAR results from the main text.